How do solve this linear equation 2(x-5)=-11?
Answer by blwinbbbles(106)

Just multiply out the problem then solve for x.
2(x-5)= -11
2x-10 = -11 add 10 to both sides
2x = -1 divide by 2
x = -1/2
so the answer is
x= -1/2